OpenStudy (anonymous):

Evaluate the following: [(x+y)^2]/[4x^2-y^2] = ??? if x = 3 and y = -5d. I got the following: [(3+(-5d))^2]/[4(3)^2-(-5d)^2] => [9-30d+25d^2]/[36+25d^2]

OpenStudy (anonymous):

Does it end up with: 9-30d/36?

OpenStudy (anonymous):

Equation View: Evaluate the following if x = 3; y -5d \[\frac{ (x+y)^{2} }{ 4x ^{2}-y ^{2} }\] I got the following: \[\frac{ (3+(-5d))^{2} }{ 4(9) ^{2}-(-5d) ^{2} }\] \[\frac{ 9-30d+25d^{2} }{ 36+25d^{2} }\] I cant figure out the next step.

TheSmartOne (thesmartone):

Factoring it..

TheSmartOne (thesmartone):

For the numerator rearrange it to \(\ 25d^{2}-30d+9 \)

TheSmartOne (thesmartone):

And we use the perfect square trinomial which is \(\ A^2 - 2AB + B^2 = (A - B)^2\)

TheSmartOne (thesmartone):

Do you get that? @Sadworld @~ :P

TheSmartOne (thesmartone):

In this example we have \(\ \color{blue}{ A = 5x }\) and \(\ \color{red}{ B = 3 }\)

TheSmartOne (thesmartone):

So \(\ 25x^{2}-30x+9 = ( \color{blue}{ 5x } - \color{red}{ 3 } )^2\)

OpenStudy (anonymous):

So you mean I should factor it? My answer would bet the same as the question: (3+(-5d))^2

OpenStudy (anonymous):

I'm asked to simplify further. I'm wondering if 25d^2 cancel each other out making the it 9-30d/36

Directrix (directrix):

Looking at the denominator from the very beginning: [4x^2-y^2] = [36 - ( - 5d ) ^2 ] = [ 36 - ( + 25 d^2) ] = [ 36 - 25 d^2 }

Directrix (directrix):

I think you have the denominator as [ 36 + 25 d^2 } Not sure, but please check. @Sadworld

OpenStudy (anonymous):

I see, you're right

OpenStudy (anonymous):

What I should do next after solving it? I'm asked to simplify further.

Directrix (directrix):

That denominator will not factor over the real numbers.

Directrix (directrix):

Let me look again at the numerator.

Directrix (directrix):

Just in case, would you double-check and see if you posted the problem correctly?

Directrix (directrix):

Could you paste the problem here?

OpenStudy (anonymous):

Yeah, the problem is correct. I don't think it can be factored further. I just want to be sure. Just a 2nd copy of the problem: If x = 3 and y = -5b \[\frac{ (x+y)^{2} }{ 4x ^{2}-y ^{2} }\]

OpenStudy (anonymous):

Can you divided with 1/common deno?

OpenStudy (anonymous):

basically cross multiply?

Directrix (directrix):

y = -5b ? or y = -5d ? as stated in the original post?

OpenStudy (anonymous):

It's y = -5d; typo, sorry

Directrix (directrix):

I don't think the expression will simplify. Were you told that it would?

OpenStudy (anonymous):

Nope, got it in a problem book. Just says to simply as much as possible. Cross multiply with itself isn't a possibility?

Directrix (directrix):

Cross multiply when there is an equation. There is no equation here.

Directrix (directrix):

I am going to rework the problem from the beginning and see if I get the same answer. I will report back here on my result, okay?

OpenStudy (anonymous):

Sure thing, I'll be online for awhile. I can check up tom is it takes longer.

Directrix (directrix):

It should go fast.

OpenStudy (anonymous):

BTW, I've factored/simplified the originals question, dunno if it matters: \[\frac{ (x+y)(x+y) }{ (2x+y)(2x-y) }\]

Directrix (directrix):

No common factors upstairs and downstairs. Okay, my re-check gives Numerator: (3 - 5d)^2 Denominator: (36 - 25 d^2) = ( 6 + 5d) (6 - 5d)

OpenStudy (anonymous):

same here

Directrix (directrix):

I think that is the answer. Usually, these books read "simplify if possible" so as not to make us think the result should simplify.

OpenStudy (anonymous):

Okay, thanks. So I've been knocking on my head for no reason. Well better try then be sorry :)

Directrix (directrix):

The "cleanest" way to write this,I think, is with Numerator: (3 - 5d)^2 Denominator: (36 - 25 d^2)

Directrix (directrix):

Right, I agree. And, we did find a sign error even if it did not allow us to strike out factors.

Directrix (directrix):

Hey, if this turns out to be a different answer, would you post here and let me know. I just cannot imagine what a different answer would be.

OpenStudy (anonymous):

Sure thing, if there's no answer then we got it right :D

Directrix (directrix):

I'm laughing at that. And, I enjoyed exchanging math ideas with you. Let's do it again sometime, okay?